OpenStudy (warpedkitten):

What is the multiplicative identity of the complex number –3 + 5i? 0 1 –3 + 5i 3 – 5i

OpenStudy (anonymous):

Do you know what multiplicative identity means?

OpenStudy (warpedkitten):

Well the answer was B. 1. If I knew what it was I wouldn't have asked for assistance..

OpenStudy (anonymous):

Well then, how did you get the answer? I'm only trying to help.

OpenStudy (anonymous):

lets call the multiplicative identity a then it must be that (–3 + 5i)a=a(–3 + 5i)=(–3 + 5i) (3 – 5i)(3 – 5i)=9-30i-25=-16-30i what number when you multiply another number by it you get that other number back? 6x=6 what is x? its 1

OpenStudy (warpedkitten):

I guessed on the answer and got it right. I didn't know how to get the answer.